OpenStudy (anonymous):

Provide a unique real world example where quadratic equations might be used.

OpenStudy (anonymous):

i dont know how i would make this into a real world situation

Directrix (directrix):

Ray operates a small sign-making business. He finds that if he charges x dollars for each sign he sells 40−x signs per week. What is the least number of signs he can sell to have an income of $336 in one week?

OpenStudy (experimentx):

good question, when a body is falling from height h, them height can be modeled on quadratic equation -h + at + bt^2 = 0

OpenStudy (experimentx):

when you shoot a projectile, the path traced can be modelled with a quadratic equation - y + bx + cx^2 = 0

OpenStudy (experimentx):

anything that looks like parabolic or parabola ... can be modeled on quadratic equation

OpenStudy (anonymous):

well, the orbit of the earth around the sun is elliptical.... and we can write a quadratic equation for that orbit which is an equation for an ellipse given by, \[(x^2/a^2)+(y^2/b^2)=1\]

Directrix (directrix):

@felixsalter --> do you need the solution?